OpenStudy (anonymous):

find (f-1)'(x) for f(x)=cos2x on the interval of (0,pi/2)

OpenStudy (anonymous):

I have already found f(-1)'(x)=\[\frac{ -1 }{ 2\sqrt{1-x^2} }\]

OpenStudy (anonymous):

yeah the derivative of the inverse. This problem I have though is, what am I supposed to do with the interval?

OpenStudy (callisto):

Originally, we have domain = (0,pi/2) , range = ...?

OpenStudy (anonymous):

doesn't say, the only thing they give me is the function and on 0,pi/2

OpenStudy (anonymous):

oh wait sorry, we can find the range. The range is 1,-1?

OpenStudy (callisto):

(1, -1), the brackets are important. For inverse function, the range will be the domain..

OpenStudy (anonymous):

ok but im still not sure what I have to be answering here. Am I supposed to plug in the values for x and find the difference like integrals?

OpenStudy (callisto):

No, I don't think you need to..

OpenStudy (anonymous):

so why would they give me the interval? do I just find the inverse of the derivative?

OpenStudy (callisto):

Not quite sure... @hartnn Please give a hand here!!

OpenStudy (anonymous):

so would this be all?

OpenStudy (anonymous):

ignore that part of the question

OpenStudy (anonymous):

the words "on the interval of \([0,\frac{\pi}{2}]\)" are necessary to make the function \(\cos(2x)\) a one to one function if the interval were larger, it would not be one to one and therefore would not have an inverse

OpenStudy (anonymous):

it is like saying "let \(f(x)=\frac{1}{x-1}, x\neq 1\)" just restricting the domain to make the function well defined

OpenStudy (anonymous):

ah ok, so my final answer is correct then?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh, maybe hold on before i say something stupid

OpenStudy (anonymous):

yeah you are right, sorry about that

OpenStudy (anonymous):

haha ok thanks so much! It always messes me up when they add something that im not going to use

OpenStudy (anonymous):

yeah i know often people think the extra restrictions are part of the question, when it is really part of the definition of the function a result of bad notation really

OpenStudy (anonymous):

haha ok thanks again! and thanks for your help too @Callisto !

OpenStudy (callisto):

Thanks @satellite73 and thanks for your question as well - good revision :)